OpenStudy (zzelinski):

The sum of three numbers is 10. The first number minus the second number plus the third is 6. The first minus the third is 2 more than the second. Find the numbers. I've got so far, x + y + z =10 x - y + 3 = 6 I can't figure out the third equation.

OpenStudy (anonymous):

x-z=2+y

OpenStudy (anonymous):

I think

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

now solve the ssytem

OpenStudy (zzr0ck3r):

x + y + z =10 x - y + 3 = 6 x-z=2+y solve the third for x x= 2+y+z plug that into the second and solve for y 2+y+z - y + 3 = 6 z+5 = 6 z=1 x+y+z=10 x+y+1=10 2+y+1+y+1=10 2y+4 = 10 y=3 x = 6

OpenStudy (zzelinski):

Thanks so much for your help!

OpenStudy (whpalmer4):

@zzr0ck3r Hold on, I think you did the wrong equations! The sum of three numbers is 10. x+y+z = 10 The first number minus the second number plus the third is 6. x-y+z = 6 The first minus the third is 2 more than the second. Find the numbers. x-z=y+2

OpenStudy (whpalmer4):

Different answer as a result...

OpenStudy (zzelinski):

This isn't right....because 6 + 3 + 1 equals 10 but 6 - 3 + 1 doesn't equal 6...

OpenStudy (zzr0ck3r):

woops, sorry

OpenStudy (whpalmer4):

yeah, you read "third" as 3 instead of "z"...

OpenStudy (zzr0ck3r):

yeah, someone fix:) I got to go take a final :( sorry

OpenStudy (whpalmer4):

This illustrates my argument that you *always* need to plug the answers back into all of the equations and make sure they actually are a valid solution :-)

OpenStudy (whpalmer4):

So, thanks! Good luck with the final :-)

OpenStudy (whpalmer4):

@zzelinski do you want to try solving the corrected system?

OpenStudy (zzelinski):

Yeah I was checking it and it didn't make sense...and yes I would!

OpenStudy (whpalmer4):

(medal awarded for checking and spotting the error!)

OpenStudy (zzelinski):

@zzr0ck3r thanks for the help anyways and good luck on your final!

OpenStudy (zzr0ck3r):

I think I checked the first one and thought it was good. x+y+z=10 x-y+z=6 x-z=2+y so x= 2+y+z so x-y+z=6 ----> 2+y+z-y+z=6 2z=4 z=2 x+y+z=10 ---> 2+y+2+y+2=10 2y=4 y=2 x = 2+2+2 = 6

OpenStudy (zzr0ck3r):

x=6, y=2, z=2

OpenStudy (zzr0ck3r):

I think that's right:)

OpenStudy (zzr0ck3r):

I hope you see how I did it, if not ask any questions.

OpenStudy (whpalmer4):

Yeah, with a system of equations, it's possible to have solutions to some equations turn out correct even though others do not. that's why you need to try them all. last set of answers agrees with mine, and check out: 6+2+2 = 10, 6-2+2=6, 6-2=2+2 :-)

OpenStudy (zzelinski):

Yes! Thank you all so much! I worked through the problem and got (6,2,2) as well.

OpenStudy (whpalmer4):

Great! When checking my work, I like to reread the original statement, too — I discovered the first time I worked this problem that I'd thought the second sentence was "the first number minus 4 times the second number ..." and had I simply checked my solution in my equations, I wouldn't have noticed the error.

OpenStudy (zzr0ck3r):

Or read the thing at all... I just went with what they said the equations was:)